# MoCE: Adaptive Mixture of Contextualization Experts for Byte-based Neural Machine Translation

### v1

@@ Abstract
Byte-based machine translation systems have shown significant potential in massively multilingual settings. Unicode encoding, which maps each character to specific byte(s), eliminates the emergence of unknown words, even in new languages, enabling broad language scalability. However, byte-level tokenization results in sequences that are hard to interpret due to limited semantic information per byte. Local contextualization has proven effective in assigning initial semantics to tokens, improving sentence comprehension. Nevertheless, variations in encoding rules across languages necessitate an adaptive approach for effective contextualization. To this end, we propose Adaptive MultiScale-Headed Attention (Ada-MSHA), adaptively selecting and mixing attention heads, which are treated as contextualization experts. This enhances the flexibility of contextualization scales and improves the potential to discover a better strategy than previous methods. Experiment results show that our method outperforms existing methods without extensive manual adjustment of hyper-parameters and surpasses subword-based models with fewer parameters in Ted-59 dataset. Our code is available at https://github.com/ictnlp/MoCE.
@@
Langlin Huang 1,3, Mengyu Bu 1,3, Yang Feng 1,2,3††thanks: Corresponding author. 1 Key Laboratory of Intelligent Information Processing Institute of Computing Technology, Chinese Academy of Sciences 2 Key Laboratory of AI Safety, Chinese Academy of Sciences 3 University of Chinese Academy of Sciences h.langlin@wustl.edu, {bumengyu23z, fengyang}@ict.ac.cn
@@ 1 Introduction
Neural Machine Translation (NMT) is a consistently hot research topic, and recent years have seen the growing significance of multilingual language modeling Zhang et al. (2023). The selection of tokenization and vocabulary is critical to multilingual language models, which plays an important role in vectorization of texts and discretization of predicted hidden states. While some models Costa-jussà et al. (2022); Dubey et al. (2024) use large vocabularies to ensure word coverage, others Touvron et al. (2023); Jiang et al. (2023) opt for byte fallback strategy. This approach allows them to completely avoid unknown words with a smaller vocabulary size. Byte-based models like ByT5 Xue et al. (2022), MEGABYTE Yu et al. (2023) and Shaham and Levy (2021) convert all words into UTF-8 byte, which further reduces the vocabulary size to about 256. This strategy significantly reduces the size of the embedding table, saving parameters and accelerating the vectorization and discretization. Besides, it eliminates the unknown-word problem and can be easily generalized to massively multilingual scenarios. Empirical study Edman et al. (2024) has also shown the performance superiority of byte-based MNT models.
@@
Although MSC provides an effective framework for modeling multi-scale contextualization and achieved state-of-the-art performance, it suffers from a significant limitation: the scales are manually predefined. This reduces the model’s ability to generalize to multilingual scenarios, particularly in massively multilingual machine translation, which may involve over 50 languages. Under UTF-8 rule, a character may convert to 1 to 4 bytes, depending on the language. This leads to varying requirements of contextualization scale for different languages. However, once MSC decides the contextualization scales, they are unchangeable for any input. To address this, we leverage the concept of Mixture of Experts (MoE) Shazeer et al. (2017) and propose Mixture of Contextualization Experts (MoCE), which can adaptively determine CNN kernel sizes based on each input text.
Specifically, we modify Multi-Head Attention (MHA) to propose a structure, called Adaptive MultiScale-Headed Attention (Ada-MSHA). This attention layer allows each head to be locally contextualized and the contextualization scales are adaptive to the input. Instead of predefined scales that MSC uses, Ada-MSHA dynamically combines different scales with model needs. The flexibility of contextualization scales is therefore significantly enhanced, resulting in better performance. Additionally, we prove that given language ID as prior knowledge is beneficial to the selection of scales.
Experiment results on two massively multilingual translation datasets, Ted-59 and OPUS-100, demonstrate our proposed method outperforms other byte-based translation models with very few extra parameters. Comparing with the subword-based model, Ada-MSHA requires fewer parameters while performing better in Ted-59 dataset.
@@ 2.1 Mixture-of-Experts (MoE)
MoE was designed mainly to increase the potential parameters of a model Shazeer et al. (2017); Lepikhin et al. (2021); Jiang et al. (2024). Recently, it is also used in multi-domain Du et al. (2023) or multi-task Park (2024) scenarios.
@@ 3 Method
Leveraging multi-scale contextualization, MSC Huang and Feng (2024) has achieved the state-of-the-art performance on byte-based NMT. However, it has a critical drawback that the contextualization scales are predefined and fixed. Worse still, it costs many trials to figure out a suitable set of these hyper-parameters, which can change largely among different datasets. In this section, we propose Adaptive MultiScale-Headed Attention (Ada-MSHA) that can adjust to not only different datasets but also various languages.
@@ 3.2 Adaptive MultiScale-Headed Attention
It should be clarified that Ada-MSHA is a modified Attention layer. By default, we replace the first encoder layer with Ada-MSHA and keep the rest model parts unchanged. This is discussed in Appendix F.
@@ 3.3 Language ID - the Prior Information
The experiment results in H demonstrate the "" is beneficial to massively multilingual translation.
@@ 4.2 Systems
Ada-MSHA: Our proposed method, with hyper-parameter or . Routing with language ID is also tested, shown as "".
@@ 4.3 Training, Inference, and Evaluation
We follow the standard practice in multilingual translation by training models on both "xxen" and "enxx" translation directions. Since the focus is handling diverse source languages, our evaluation primarily targets the "xxen" direction. Apart from reporting BLEU scores in the main body, we also report character-level metric ChrF and model-based metric COMET in Appendix D and C for comprehensive evaluation. We detail the complete setups in Appendix B.
@@ 5.3 Main Experiments
Table 3 summarizes the results on Ted-59 dataset. Nearly all byte-based models outperform the subword-based model with significantly fewer parameters. Among the byte-based models, all varieties of Ada-MSHA surpass MSC by a large margin, demonstrating the effectiveness of using adaptive contextualization scales.
The comparison of different settings of Ada-MSHA demonstrates two points. First, applying "" achieves consistent improvement, which is discussed in Appendix H. Second, value affects model’s inclination toward language groups, especially "Long" and "Medium" groups, which is discussed in Section 5.5. The results at "Short" group do not follow a certain rule. We conjecture this is because a single byte has a determined mapping to character for "Short" languages, so they rely less on contextualization functions.
@@ 5.4 Expert Choice and Language Conciseness
The major motivation of Ada-MSHA is to choose appropriate contextualization functions according to the input. To verify if Ada-MSHA achieves such target, we count the selected ratio of each expert throughout a whole test set. Specifically, we select 4 languages, my, bg, et, and zh, from left to right in Figure 3, and also from least concise to most concise. Then, we test "enxx" and "xxen" for them, recording the selected ratios of experts. The sentence length ratios are recorded to represent the relative conciseness, which highly correlates to the expert choices. All experiments are conducted on Ted-59 with "+" setting.
@@
These findings demonstrate that Ada-MSHA can effectively identify the input language type and route the input to appropriate experts.
@@ 5.5 Systematic Shift of with Varied
Comparing and in Figure 4, we observe a significant shift of selection. This shift of overall tendency is systematic, caused by model structure change. The vertical lines quantitatively demonstrate a model with larger tends to select experts with a larger radius. This explains why performs better in "Long" and performs better in "Medium".
@@ 5.7 Effectiveness of MoCE Concept
We highlight the benefit of applying a Mixture of Contextualization Experts over using a single expert for each head. With the weights of experts varying in continuous space, the defacto contextualization radius has way more possibilities. Here, we choose top-1 routing to simulate the single expert setting and compare it with Ada-MSHA. Empirical results in Table 5 demonstrate this advantage.
@@ 5.8 Training and Inference Speed
While Ada-MSHA demonstrates substantial improvements in translation quality over the baselines, the concern of time cost may rise. To address this concern, we evaluated the training and inference speed of all methods mentioned in this paper using Ted-59 dataset, providing a quantitative comparison of computational efficiency. The byte-based Transformer model serves as the baseline, the relative speed ratios are exhibited in Table 6.
@@
For the training phase, we recorded the time spent in training one epoch. For the inference phase, we recorded the time spent in translating the entire test set for each translation direction, and took the sum of them. The speed of Transformer (byte) is denoted as "1x"; values below 1 indicate slower speeds compared to the baseline. According to Table 6, all the local contextualization methods, including Ada-MSHA, achieve nearly the same speed as the baseline.
@@ 6.3 Local Contextualization Modeling
While the attention mechanism’s global contextualization effectively addresses the challenge of long-distance dependencies, many tasks benefit from a focus on local contextualization. For instance, in audio processing, a model needs several consecutive frames to comprehend an audio segment. The Conformer Gulati et al. (2020) has proven to be an effective local contextualization structure for audio-related tasks such as speech recognition Gulati et al. (2020), speech translation Fang et al. (2024), and simultaneous translation Ma et al. (2024). Given that Ada-MSHA provides similar functionality to Conformer, it may serve as a more advantageous replacement. Moreover, tasks involving phrasal representation modeling Huang et al. (2023); Fang and Feng (2022) could also benefit from Ada-MSHA’s capacity as a robust multi-scale feature extractor.
@@ Appendix B Detailed Training, Inference and Evaluation Setup
For comprehensive evaluation, we report BLEU Papineni et al. (2002) and ChrF Popović (2015) scores using SacreBLEU toolkit Post (2018), and report COMET Rei et al. (2022) score using wmt22-comet-da555https://huggingface.co/Unbabel/wmt22-comet-da model. To keep the article concise, we present BLEU scores in the main body, and present ChrF and COMET in the Appendix C and D, respectively.
Appendix C ChrF Scores on Ted-59 and OPUS-100
@@
Appendix D Comet Scores on Ted-59 and OPUS-100
@@
Appendix E Experiment on a Small Multilingual Translation Dataset
@@
Appendix F The Best Position to Place Ada-MSHA Layer
@@
Appendix G Experiment Under Low Resource Setting
@@
The training, inference, and evaluation details are almost the same as those for Ted-59 dataset. The only difference is the representative languages for the Long, Medium, and Short categories. Since some of the languages from Table 2 are not low resources, we re-selected four languages for each category, the first two from L.R. and the last two from E.L.R., according to the order in Figure 3. For Ada-MSHA, we experiment only with the best model setting, with "lid".
The experiment results are exhibited in Table 12. Surprisingly, byte-based Transformer fails the tasks, while subword-based Transformer performs adequately. This demonstrates the difficulty of learning byte-based text with only global attention. Compared with byte-based Transformer, other local contextualization-based methods are able to learn the patterns with the low-resourced training data. Among them, Ada-MSHA shows significant improvement over other methods, and the advantages are larger than that on the full Ted-59 dataset. This demonstrates our proposed method, which adaptively chooses contextualization neighbors, requires fewer data to converge to an optimal solution.
@@
Appendix H Effectiveness of Using Language ID

### v2

@@ Abstract
Byte-based machine translation systems have shown significant potential in massively multilingual settings. Unicode encoding, which maps each character to specific byte(s), eliminates the emergence of unknown words, even in new languages. This avoids out-of-vocabulary risk in multilingual translation and enables broad language scalability. However, byte-level tokenization results in sequences that are hard to interpret due to limited semantic information per byte. Local contextualization has proven effective in assigning initial semantics to tokens, improving sentence comprehension. Nevertheless, variations in encoding rules across languages necessitate an adaptive approach for effective contextualization. To this end, we propose Mixture of Contextualization Experts (MoCE), adaptively selecting and mixing attention heads, which are treated as contextualization experts. This enhances the flexibility of contextualization scales and allows models to search for better contextualization combinations. Experiment results show that our method outperforms existing methods without extensive manual adjustment of hyper-parameters and surpasses subword-based models with fewer parameters in Ted-59 dataset. Our code is available at https://github.com/ictnlp/MoCE.
@@
Langlin Huang 1,3,4222This paper was done when Langlin Huang studied at Institute of Computing Technology, Chinese Academy of Sciences., Mengyu Bu 1,3, Yang Feng 1,2,3111Corresponding author. 1 Key Laboratory of Intelligent Information Processing Institute of Computing Technology, Chinese Academy of Sciences 2 Key Laboratory of AI Safety, Chinese Academy of Sciences 3 University of Chinese Academy of Sciences 4 Washington University in St. Louis h.langlin@wustl.edu, {bumengyu23z, fengyang}@ict.ac.cn
@@ 1 Introduction
Neural Machine Translation (NMT) is a consistently hot research topic, and recent years have seen the growing significance of multilingual language modeling Zhang et al. (2023). The selection of tokenization and vocabulary is critical to multilingual language models, which plays an important role in vectorization of texts and discretization of predicted hidden states. While some models Costa-jussà et al. (2022); Dubey et al. (2024) use large vocabularies to ensure word coverage, others Touvron et al. (2023); Jiang et al. (2023) opt for byte fallback strategy. These approaches allow them to completely avoid unknown words with a smaller vocabulary size. Byte-based models like Xue et al. (2022); Yu et al. (2023); Shaham and Levy (2021) convert all words into UTF-8 byte, which further reduces the vocabulary size to about 256. This strategy also reduces the size of the embedding table, saving parameters and accelerating token embedding and inference. Besides, it eliminates the unknown-word problem and can be easily generalized to massively multilingual scenarios. Empirical study Edman et al. (2024) has also shown the performance superiority of byte-based MNT models.
@@
Although MSC provides an effective framework for modeling multi-scale contextualization and achieves state-of-the-art performance, it suffers from the limitation of manually predefined scales. This reduces the model’s ability to generalize to multilingual scenarios, particularly in massively multilingual machine translation, which may involve over 50 languages. Under UTF-8 rule, a character may convert to 1 to 4 bytes, depending on the language. This leads to varying requirements of contextualization scale for different languages. However, once MSC decides the contextualization scales, they are unchangeable for any input.
To address this, we leverage the concept of Mixture of Experts (MoE) Shazeer et al. (2017) and propose Mixture of Contextualization Experts (MoCE), which can adaptively determine CNN kernel sizes based on each input text. Specifically, we modify Multi-Head Attention to propose Adaptive MultiScale-Headed Attention (Ada-MSHA) module. This proposed attention allows each head to be locally contextualized and the contextualization scales are adaptive to the input. Instead of predefined scales that MSC uses, MoCE dynamically combines different scales with model needs. The flexibility of contextualization scales is therefore significantly enhanced, resulting in better performance. Additionally, we prove that given language ID as prior knowledge benefits the scale selection.
Experiment results on two massively multilingual translation datasets, Ted-59 and OPUS-100, demonstrate our proposed method outperforms other byte-based translation models with similar parameter use. Compared with the subword-based model, MoCE requires fewer parameters while performing better in Ted-59 dataset.
@@ 2.1 Mixture-of-Experts (MoE)
MoE was designed mainly to increase the potential parameters of a model Shazeer et al. (2017); Lepikhin et al. (2021); Jiang et al. (2024). Recently, it is also used in multi-domain Du et al. (2023) or multi-task Park (2024); Huang et al. (2023a) scenarios.
@@ 3 Method
In this section, we first introduce MultiScale-Headed Attention, a way to model multi-scale contextualization. Then we extend it to Adaptive MultiScale-Headed Attention, the core module of MoCE. We also propose how to leverage language ID information to further improve our method.
@@ 3.2 Adaptive MultiScale-Headed Attention
It should be clarified that Ada-MSHA is a modified Attention layer. By default, we replace the first encoder layer with Ada-MSHA and keep the rest model parts unchanged. This is discussed in Appendix G.
@@ 3.3 Language ID - the Prior Information
The experiment results in I demonstrate the "" is beneficial to massively multilingual translation.
@@ 4.2 Systems
MoCE: Our proposed method, with hyper-parameter or . Routing with language ID is also tested, shown as "".
@@ 4.3 Training, Inference, and Evaluation
We follow the standard practice in multilingual translation by training models on both "xxen" and "enxx" translation directions. Since the focus is handling diverse source languages, our evaluation primarily targets the "xxen" direction. Apart from reporting BLEU scores in the main body, we also report character-level metric ChrF and model-based metric COMET in Appendix E and D for comprehensive evaluation. We detail the complete setups in Appendix B.
@@ 5.3 Main Experiments
Table 3 summarizes the results on Ted-59 dataset. Nearly all byte-based models outperform the subword-based model with much fewer parameters. Among the byte-based models, all varieties of MoCE surpass MSC by a large margin, demonstrating the effectiveness of using adaptive contextualization scales.
The comparison of different settings of MoCE demonstrates two points. First, applying "" achieves consistent improvement, which is discussed in Appendix I. Second, value affects model’s inclination toward language groups, especially "Long" and "Medium" groups, which is discussed in Section 5.5. The results at "Short" group do not follow a certain rule. We conjecture this is because a single byte has a determined mapping to character for "Short" languages, so they rely less on contextualization functions.
@@ 5.4 Expert Choice and Language Conciseness
The major motivation of MoCE is to choose appropriate contextualization functions according to the input. To verify if MoCE achieves such target, we count the selected ratio of each expert throughout a whole test set. Specifically, we select 4 languages, my, bg, et, and zh, from left to right in Figure 3, and also from least concise to most concise. Then, we test "enxx" and "xxen" for them, recording the selected ratios of experts. The sentence length ratios are recorded to represent the relative conciseness, which highly correlates to the expert choices. All experiments are conducted on Ted-59 with "+" setting.
@@
These findings demonstrate that MoCE can effectively identify the input language type and route the input to appropriate experts.
@@ 5.5 Systematic Shift of with Varied
Comparing and in Figure 4, we observe a consistent shift of selection. This shift of overall tendency is systematic, caused by model structure change. The vertical lines quantitatively demonstrate a model with larger tends to select experts with a larger radius. This explains why performs better in "Long" and performs better in "Medium".
@@ 5.7 Effectiveness of MoCE Concept
We highlight the benefit of applying a Mixture of Contextualization Experts over using a single expert for each head. With the weights of experts varying in continuous space, the defacto contextualization radius has way more possibilities. Here, top-1 routing is chosen to simulate the single expert setting. Empirical results in Table 5 demonstrate the advantage of applying MoE in our method. It also proves that mixing two experts is enough.
@@ 5.8 Training and Inference Speed
While MoCE demonstrates substantial improvements in translation quality over the baselines, the concern of time cost may rise. To address this concern, we evaluated the training and inference speed of all methods mentioned in this paper using Ted-59 dataset, providing a quantitative comparison of computational efficiency. The byte-based Transformer model serves as the baseline, the relative speed ratios are exhibited in Table 6.
@@
For the training phase, we recorded the time spent in training one epoch. For the inference phase, we recorded the time spent in translating the entire test set for each translation direction, and took the sum of them. The speed of Transformer (byte) is denoted as "1x"; values below 1 indicate slower speeds compared to the baseline. According to Table 6, all the local contextualization methods, including MoCE, achieve nearly the same speed as the baseline.
@@ 6.3 Local Contextualization Modeling
While the attention mechanism’s global contextualization effectively addresses the challenge of long-distance dependencies, many tasks benefit from a focus on local contextualization. For instance, in audio processing, a model needs several consecutive frames to comprehend an audio segment. The Conformer Gulati et al. (2020) has proven to be an effective local contextualization structure for audio-related tasks such as speech recognition Gulati et al. (2020), speech translation Fang et al. (2024), and simultaneous translation Ma et al. (2024). Given that MoCE provides similar functionality to Conformer, it may serve as a more advantageous replacement. Moreover, tasks involving phrasal representation modeling Huang et al. (2023b); Fang and Feng (2022) could also benefit from MoCE’s capacity as a robust multi-scale feature extractor.
@@ Appendix B Detailed Training, Inference and Evaluation Setup
For comprehensive evaluation, we report BLEU Papineni et al. (2002) and ChrF Popović (2015) scores using SacreBLEU toolkit Post (2018), and report COMET Rei et al. (2022) score using wmt22-comet-da555https://huggingface.co/Unbabel/wmt22-comet-da model. To keep the article concise, we present BLEU scores in the main body, and present ChrF and COMET in the Appendix D and E, respectively.
Appendix C Applying MoCE on Subword-based Models
The experiment results in Section 5 have demonstrated the significant effectiveness of MoCE on byte-based models. In addition to that, we also explored how it benefits subword-based models. Intuitively, subword tokens contain richer information and require less contextual information. Therefore, we experimented on shorter contextual scales, i.e. and . We report the BLEU scores on Table 7. According to the results, applying MoCE improves subword-based models, though the improvement is relatively minor. We conjecture this is because the subword-based vocabulary still suffers from the curse of multilinguality.
| Long | Medium | Short | L.R. | H.R. | All | |
| Transformer | 15.43 | 31.05 | 26.34 | 23.57 | 30.69 | 24.79 |
| MoCE () | 15.71 | 31.15 | 26.65 | 23.49 | 31.05 | 24.99 |
| +lid | 15.51 | 31.46 | 26.66 | 23.77 | 31.10 | 24.98 |
[…]
| +lid | 15.63 | 31.46 | 27.26 | 24.26 | 31.59 | 25.34 |
Appendix D ChrF Scores on Ted-59 and OPUS-100
@@
Appendix E Comet Scores on Ted-59 and OPUS-100
@@
Appendix F Experiment on a Small Multilingual Translation Dataset
@@
Appendix G The Best Position to Place Ada-MSHA Layer
@@
Appendix H Experiment Under Low Resource Setting
@@
The training, inference, and evaluation details are almost the same as those for Ted-59 dataset. The only difference is the representative languages for the Long, Medium, and Short categories. Since some of the languages from Table 2 are not low resources, we re-selected four languages for each category, the first two from L.R. and the last two from E.L.R., according to the order in Figure 3. For MoCE, we experiment only with the best model setting, with "lid".
The experiment results are exhibited in Table 13. Surprisingly, byte-based Transformer fails the tasks, while subword-based Transformer performs adequately. This demonstrates the difficulty of learning byte-based text with only global attention. Compared with byte-based Transformer, other local contextualization-based methods are able to learn the patterns with the low-resourced training data. Among them, MoCE shows significant improvement over other methods, and the advantages are larger than that on the full Ted-59 dataset. This demonstrates our proposed method, which adaptively chooses contextualization neighbors, requires fewer data to converge to an optimal solution.
@@
Appendix I Effectiveness of Using Language ID